OpenStudy (anonymous):

Please help ! i really need this ..Fan AND medal Chris withdraws $7.50 per week from his savings account. After 15 weeks, the balance in the account is $585. A. Write an equation in point -slope form to model this. B. Write the equation in slope-intercept form. C. How much is in the account after 1 year (52 weeks

OpenStudy (anonymous):

a) y2-7.50=7.50x(x2-1) b) y=7.50+697.5 c) rounded to the nearest tenth the answer is 15.8